The distribution of grades in an introductory finance class is normally distributed, with an expected grade of 65. If the standard deviation of grades is 14, in what range would you expect 90.00 percent of the grades to fall? (Round answers to 2 decimal places, e.g. 15.25. Hint: Think in terms of what the expected highest and lowest scores would be for 90.00% of the students taking the exam.)

90.00 percent of the grades should fall beween the ranges (lower end)

and (upper end)
Answer #1

margin of error = z-score * standard deviation

= 1.645 * 14

= 23.03

lower bound = 65 - 23.03

= 41.97

upper bound = 65 + 23.03

= 88.03

hence range is 41.97 to 88.03